### v1

# Inferring neutron-star Love-Q relations from gravitational waves in the hierarchical Bayesian framework

Abstract
Despite the large uncertainties in the equation of state for neutron stars (NSs), a tight universal “Love-Q” relation exists between their dimensionless tidal deformability, $\Lambda$, and the dimensionless quadrupole moment, $Q$. However, this relation has not yet been directly measured through observations. Gravitational waves (GWs) emitted from binary NS (BNS) coalescences provide an avenue for such a measurement. In this study, we adopt a hierarchical Bayesian framework and combine multiple simulated GW events to measure the Love-Q relation. We simulate 1000 GW sources and select 20 events with the highest signal-to-noise ratios and NS spins for the analysis. By inspecting four parameterization models of the Love-Q relation, we observe strong correlations between the model parameters. We verify that a linear relation between $\ln\Lambda$ and $\ln Q$ is practically sufficient to describe the Love-Q relation with the precision expected from next-generation GW detectors. Furthermore, we utilize the inferred Love-Q relation to test modified gravity. Taking the dynamical Chern-Simons gravity as an example, our results suggest that the characteristic length can be constrained to $10\,\mathrm{km}$ or less with future GW observations.
1 Introduction
Thanks to their extreme densities and strong gravitational fields, neutron stars (NSs) serve as natural laboratories for studying nuclear and gravitational physics (see e.g., Ref. [1]). Inferred quantities from electromagnetic observations, such as the observed maximum mass [2, 3, 4] and mass-radius relation [5, 6, 7, 8, 9], allow one to probe the properties of nuclear matter at densities exceeding the nuclear saturation density. Additionally, the observation of GW170817 has opened up a new window for investigating NS properties using gravitational waves (GWs) emitted from binary NS (BNS) coalescences [10, 11, 12]. Some of the NS properties, including the tidal deformability and the spin-induced quadrupole moment, contribute to the GW emission [13, 14, 15, 16, 17, 18] and therefore could be measured from GW observations [19, 20, 21, 22, 23, 24, 25].
In a BNS system, each NS is deformed due to the gravitational field of its companion, leading to an induced mass quadrupole moment [26, 27]. The effect is characterized by the tidal deformability $\Lambda=2k_{2}/(3C^{5})$, where $k_{2}$ is the tidal Love number and $C$ is the NS compactness [28]. Also, a rotating NS experiences another deformation due to its spin, which induces the so-called spin-induced quadrupole moment, $\mathcal{Q}=-Q\chi^{2}m^{3}$, where $m$ is the mass of the NS, $\chi$ and $Q$ are the dimensionless spin and quadrupole moment respectively [29, 30]. These properties can provide an insight into the internal structure of NSs and bring opportunities to test strong-field gravity [31, 32, 33, 34, 35, 36, 37].
The internal structure of NSs depends on both the underlying gravity theory and the equation of state (EOS); the latter describes the relation between pressure and density of the NS matter. Despite starting to place constraints on the EOS, current observations are in general not accurate enough to distinguish between various EOS candidates [5, 6, 7, 3, 8]. This means that when it comes to testing strong-field gravity, one will encounter a degeneracy between the EOS and the underlying gravity theory [38, 39, 40, 41, 42, 1]. One way to break this degeneracy is to find universal relations among NS properties. Yagi and Yunes [38, 39] found such a universal relation between $I$ (moment of inertia), $\Lambda$ (tidal Love number) and $Q$ (quadrupole moment), assuming the validity of general relativity (GR). On the one hand, this “I-Love-Q” relation is insensitive to EOS uncertainties with variations of about $1\%$ or less for different EOSs. On the other hand, the relation can deviate from the GR prediction in modified gravity theories [43, 44, 45], enabling an EOS-insensitive test of gravity. Among the I-Love-Q trio, the Love-Q relation can be probed with GW observations, since both $\Lambda$ and $Q$ affect the GWs emitted by BNS systems. If the Love-Q relation in GR is adopted as a prior in the waveform model, the number of independent parameters can be reduced, which helps to better estimate the spin parameters of BNSs [38, 11, 12, 46]. The existence of such universal relations, which now include many others [47, 48, 49, 50, 51, 52, 53], may also imply a “no-hair-theorem-like” behavior for NSs, bringing us a new insight into fundamental physics.
Future next-generation (XG) ground-based GW detectors, including the Cosmic Explorer (CE) [54, 55] and the Einstein Telescope (ET) [56, 57, 58, 59], are expected to detect many more GW signals, up to about $10^{5}$$10^{6}$ events per year [60, 61, 62, 63], thanks to their increased sensitivity and lower cutoff frequencies. These high-precision GW observations allow us to treat $\Lambda$ and $Q$ as independent parameters in the waveform model, to be measured directly from GWs. This enables further constraints on the Love-Q relation from an observational perspective. Samajdar and Dietrich [64] have first performed an analysis discussing the prospects of constraining Love-Q relation with GW observations, where a weighted linear regression was performed. While the results are very valuable, as we will show below, such a treatment might miss possible degeneracy and non-Gaussianity in the posteriors of $\Lambda$ and $Q$.
For the first time, our study adopts the hierarchical Bayesian framework to infer the Love-Q relation with XG GW observations. This framework has been successfully applied in population studies of compact binary coalescences, and in probing the EOS of NSs [65, 66, 67, 68, 69, 70, 71, 72]. Regarding the fitting parameters in the Love-Q relation as hyperparameters, the hierarchical Bayesian framework separates the inferences of these hyperparameters and the single-event parameters into two layers to avoid a direct, high-dimensional inference for all unknown parameters. It significantly reduces the computational cost in combining information from multiple events. Also, the construction of the quasi-likelihood function in this framework incorporates the full shape of the posterior in single-event inference beyond Gaussianity, thus utilizes the information contained therein in a more comprehensive way.
In our implementation, we simulate 1000 GW events based on population models of NSs [73, 74, 64] and select the 20 loudest events for analysis. We find that the primary information for constraining the Love-Q relation comes from the 10 loudest GW events, consistent with results found in previous studies [68]. In the pioneering work of Samajdar and Dietrich [64], a linear relation between $\ln\Lambda$ and $\ln Q$ is adopted. By further considering four polynomial models from linear to quartic terms in fitting the relation, we quantitatively show that the linear relation is accurate enough when constraining the Love-Q relation with GWs. Additionally, we apply the inferred Love-Q relation in gravity tests. Taking the dynamical Chern-Simons (dCS) gravity as an example, we find that the characteristic length $\xi_{\rm CS}^{1/4}$—with $\xi_{\rm CS}$ the theory parameter—can be limited to $\lesssim 10\,{\rm km}$ with future GW observations.
This paper is organized as follows. In section 2 we construct the hierarchical Bayesian framework and derive the posterior of the hyperparameters. The simulation procedure is explained in section 3. We present the results of our inference and discuss the differences between different Love-Q parameterization models in section 4. We compare our inference results with the predictions in the dCS gravity in section 5. Finally, we conclude in section 6.
2 Hierarchical Bayesian Inference of Love-Q Relation
2.1 Polynomial Models of Love-Q Relation
Yagi and Yunes [38, 39, 43] fit the Love-Q relation with a quadratic polynomial model as,
$\ln Q_{5}=a_{5}+b_{5}\ln\Lambda+c_{5}\ln^{2}\Lambda+d_{5}\ln^{3}\Lambda+e_{5}\ln^{4}\Lambda\,$, (2.1)
where the dimensionless fitting coefficients are $a_{5}=0.1940$, $b_{5}=0.09163$, $c_{5}=0.04812$, $d_{5}=-4.283\times 10^{-3}$ and $e_{5}=1.245\times 10^{-4}$ [43], and the lower subscript “55” indicates five model parameters. They found that such a relation applies to most of the EOSs with relative deviation less than $1\%$ in GR. When constraining the Love-Q relation with GWs, Samajdar and Dietrich [64] adopted a linear model,
$\ln Q_{2}=a_{2}+b_{2}\ln\Lambda\,$, (2.2)
where only two parameters, $a_{2}$ and $b_{2}$, are involved. In figure 1, we plot the Yagi-Yunes relation (2.1) along with its linear fit (2.2), which is obtained by a least squares regression performed on 1000 points uniformly picked in the logarithmic space from the Yagi-Yunes relation ($a_{2}=-0.1457$ and $b_{2}=0.3094$). As an example of a typical EOS, we plot the Love-Q relation calculated assuming the APR4 EOS [75], a soft EOS consistent with the observations of GW170817 [10, 11, 12]. In this case, the relative differences in $Q$ between these two models and the APR4 EOS are smaller than $1\%$.
Figure 1: Illustration of the Love-Q relation. In the upper panel, the orange solid line indicates the original Yagi-Yunes relation (2.1), while the green dashed line represents our fitting with the linear model (2.2). We also show the Love-Q relation for the APR4 EOS as reference with blue circles. In the lower panel, we show the absolute relative differences between the two models (denoted as $Q_{5}$ and $Q_{2}$) and the APR4 EOS (denoted as $Q_{\rm APR4}$).
2.2 Hierarchical Bayesian Inference
The coefficients in Eq. (2.1) and Eq. (2.2) do not directly contribute to the GW waveform. Instead, they determine a relation between the waveform parameters $\Lambda$ and $Q$, denoted as $Q=f(\Lambda;\bm{H})$, where $\bm{H}$ represents the coefficients, $\bm{H}=\{a_{2},b_{2}\}$ or $\bm{H}=\{a_{5},b_{5},c_{5},d_{5},e_{5}\}$. When the universal relation is exact, this leads to a $\delta$-function-type prior between GW parameters $\Lambda$ and $Q$
$\pi(Q|\Lambda,\bm{H})=\delta\big(Q-f(\Lambda;\bm{H})\big)\,.$ (2.3)
Intuitively, one can measure $\Lambda$ and $Q$ from BNS GW events, and then fit the relation with the measurements. This procedure can be implemented within the hierarchical Bayesian framework, which is a powerful formalism in studying population properties of GW events beyond individual observations [76].
In the population inference scenario, the population properties are characterized by a set of hyperparameters, such as the power index of the mass distribution. These hyperparameters do not enter the waveform directly, while can be inferred from a collection of measurements of single-event parameters. Similarly, the EOS parameters determine the relation between $\Lambda$ and $m$, as well as between $Q$ and $m$, can be regarded as hyperparameters and analyzed in the hierarchical Bayesian framework [65, 66, 67, 68, 69, 70, 71, 72]. In this work, we adopt this framework to the inference of the Love-Q relation.
Below we briefly introduce the hierarchical Bayesian framework and describe the customization in inferring the Love-Q relation. The hierarchical Bayesian framework aims to find the posterior distribution of the hyperparameters $p(\bm{H}|D)$, given the catalog-level data $D=\{d_{1},\cdots,d_{n}\}$ consisting of $n$ individual GW events. Since both the hyperparameters $\bm{H}$ and the single-event parameters $\{\bm{\theta}_{1},\cdots,\bm{\theta}_{n}\}$ are unknown, we write the Bayes’ theorem as
$p(\bm{H},\bm{\theta}_{1},\cdots,\bm{\theta}_{n}|D)=\frac{p(D|\bm{H},\bm{\theta}_{1},\cdots,\bm{\theta}_{n})\pi(\bm{H},\bm{\theta}_{1},\cdots,\bm{\theta}_{n})}{p(D)}\,$, (2.4)
where $\pi(\bm{H},\bm{\theta}_{1},\cdots,\bm{\theta}_{n})$ is the prior, $p(D|\bm{H},\bm{\theta}_{1},\cdots,\bm{\theta}_{n})$ is the likelihood function, and $p(D)$ denotes the evidence. Then, $p(\bm{H}|D)$ can be obtained by marginalizing over the single-event parameters
$p(\bm{H}|D)=\int p(\bm{H},\bm{\theta}_{1},\cdots,\bm{\theta}_{n}|D)\text{d}\bm{\theta}_{1}\cdots\text{d}\bm{\theta}_{n}\,.$ (2.5)
Assuming that the $n$ events are independent, the prior can be decomposed into parts of hyperparameters and that of single-event parameters,
$\pi(\bm{H},\bm{\theta}_{1},\cdots,\bm{\theta}_{n})=\pi(\bm{H})\prod_{i=1}^{n}\pi(\bm{\theta}_{i}|\bm{H})\,.$ (2.6)
Since we are inferring the Love-Q relation, the tidal deformabilities, $\bm{\Lambda}_{i}=\{\Lambda_{1i},\Lambda_{2i}\}$, and quadrupole moments, $\bm{Q}_{i}=\{Q_{1i},Q_{2i}\}$, of the two NSs of the $i\text{-th}$ event are treated as independent parameters in $\bm{\theta}_{i}$. Note that the Love-Q relation is insensitive to EOS, the priors of mass parameters, $m_{1}$ and $m_{2}$, are chosen to be independent of $\bm{\Lambda}_{i},\bm{Q}_{i}$ and $\bm{H}$. Other parameters in $\bm{\theta}_{i}$ are also assumed to be independent. In this way, the conditional prior of $\bm{\theta}_{i}$ can be further decomposed as
$\pi(\bm{\theta}_{i}|\bm{H})=\pi(\bm{\Lambda}_{i}|\bm{H})\,\pi(\bm{Q}_{i}|\bm{\Lambda}_{i},\bm{H})\,\pi(\bm{\xi}_{i})\,$, (2.7)
where $\bm{\xi}_{i}$—the so-called nuisance parameters hereafter—denotes the other parameters in $\bm{\theta}_{i}$ except for $\bm{\Lambda}_{i}$ and $\bm{Q}_{i}$. Analogously to the prior, the catalog likelihood can be factorized into the product of single-event likelihoods
$p(D|\bm{H},\bm{\theta}_{1},\cdots,\bm{\theta}_{n})=\prod_{i=1}^{n}p(d_{i}|\bm{H},\bm{\theta}_{i})=\prod_{i=1}^{n}p(d_{i}|\bm{\theta}_{i})\,$, (2.8)
where the second equality comes from the fact that the hyperparameters do not enter the waveform.
Given the specific form of the prior and the likelihood, the marginalized posterior distribution (2.5) becomes
$p(\bm{H}|D)$ $=\frac{1}{p(D)}\pi(\bm{H})\int\text{d}\bm{\theta}_{1}\cdots\text{d}\bm{\theta}_{n}\prod_{i=1}^{n}\big[\pi(\bm{\Lambda}_{i}|\bm{H})\,\pi(\bm{Q}_{i}|\bm{\Lambda}_{i},\bm{H})\,\pi(\bm{\xi}_{i})\,p(d_{i}|\bm{\theta}_{i})\big]$ (2.9)
$=\frac{1}{p(D)}\pi(\bm{H})\prod_{i=1}^{n}\int\text{d}\bm{\Lambda}_{i}\text{d}\bm{Q}_{i}\,\pi(\bm{\Lambda}_{i}|\bm{H})\,\delta\big(\bm{Q}_{i}-\bm{f}(\bm{\Lambda}_{i};\bm{H})\big)\int\text{d}\bm{\xi}_{i}\,\pi(\bm{\xi}_{i})p(d_{i}|\bm{\theta}_{i})$
$=\frac{1}{p(D)}\pi(\bm{H})\prod_{i=1}^{n}\int\text{d}\bm{\Lambda}_{i}\,\pi(\bm{\Lambda}_{i}|\bm{H})L_{i}\big(\bm{\Lambda}_{i},\bm{f}(\bm{\Lambda}_{i};\bm{H})\big)\,.$
In the second line, the bold font $\bm{f}(\bm{\Lambda}_{i};\bm{H})$ is used to denote the vector form of the Love-Q relation applied to the two NSs in a BNS system. In the third line, $L_{i}(\bm{\Lambda}_{i},\bm{Q}_{i})$, called the quasi-likelihood, is defined as the integral over the nuisance parameters
$L_{i}(\bm{\Lambda}_{i},\bm{Q}_{i})=\int\text{d}\bm{\xi}_{i}\,\pi(\bm{\xi}_{i})p(d_{i}|\bm{\theta}_{i})\,.$ (2.10)
The quasi-likelihood of each event can be computed independently of $\bm{H}$. For the $i\text{-th}$ event, we write down the Bayes’ theorem
$p(\bm{\theta}_{i}|d_{i},\varnothing)\propto\pi(\bm{\theta}_{i}|\varnothing)p(d_{i}|\bm{\theta}_{i})\,$, (2.11)
where $\pi(\bm{\theta}_{i}|\varnothing)$ denotes an auxiliary prior independent of $\bm{H}$. The explicit form of the single-event likelihood, $p(d_{i}|\bm{\theta}_{i})$, is given by assuming stationary and Gaussian noise [77]
$p(d_{i}|\bm{\theta}_{i})\propto\mathrm{e}^{-\frac{1}{2}\langle d_{i}-h(\bm{\theta}_{i}),d_{i}-h(\bm{\theta}_{i})\rangle}\,$, (2.12)
with the data $d_{i}$ and the waveform model $h(\bm{\theta}_{i})$. The inner product of $u(t)$ and $v(t)$, $\langle u,v\rangle$, is defined as
$\langle u,v\rangle:=2\int_{-\infty}^{\infty}\frac{\tilde{u}(f)\tilde{v}^{*}(f)}{S_{n}(|f|)}\text{d}f\,$, (2.13)
where $\tilde{u}(f)$ and $\tilde{v}(f)$ are the Fourier transforms of $u(t)$ and $v(t)$, and $S_{n}(f)$ is the power spectrum density (PSD) of the noise.
Rearranging terms of Eq. (2.11) and substituting it into Eq. (2.10), one finds that
$L_{i}(\bm{\Lambda}_{i},\bm{Q}_{i})=\int\text{d}\bm{\xi}_{i}\,\pi(\bm{\xi}_{i})p(d_{i}|\bm{\theta}_{i})\propto\int\text{d}\bm{\xi}_{i}\frac{\pi(\bm{\xi}_{i})}{\pi(\bm{\theta}_{i}|\varnothing)}p(\bm{\theta}_{i}|d_{i},\varnothing)\,.$ (2.14)
If we further choose $\pi(\bm{\theta}_{i}|\varnothing)\propto\pi(\bm{\xi}_{i})$, i.e., a flat prior for $\bm{\Lambda}_{i}$ and $\bm{Q}_{i}$, Eq. (2.14) can be further simplified as
$L_{i}(\bm{\Lambda}_{i},\bm{Q}_{i})\propto\int\text{d}\bm{\xi}_{i}\,p(\bm{\theta}_{i}|d_{i},\varnothing)\propto p(\bm{\Lambda}_{i},\bm{Q}_{i}|d_{i},\varnothing)\,.$ (2.15)
This means that the quasi-likelihood is proportional to the marginalized posterior of the auxiliary single-event inference with a flat prior on $\bm{\Lambda}_{i}$ and $\bm{Q}_{i}$.
The hierarchical Bayesian framework, as its name implies, introduces two levels of inferences. The first level consists of single-event Bayesian inferences based on Eq. (2.11), where the quasi-likelihoods are constructed according to Eq. (2.15). In the second level, the quasi-likelihoods are combined to infer the hyperparameters based on Eq. (2.9).
3 Simulation
3.1 Waveform, Population and Detectors
In our simulation, we adopt the IMRPhenomXAS_NRTidalv3 waveform model [18], which includes tidal amplitude corrections as well as spin-induced quadrupole moment terms up to 3.5 PN with aligned spins. The single-event parameters $\bm{\theta}$ include the binary masses $m_{1}$ and $m_{2}$, the dimensionless tidal deformabilities $\Lambda_{1}$ and $\Lambda_{2}$, spin-induced quadrupole moments $Q_{1}$ and $Q_{2}$, the dimensionless spins $\chi_{1}$ and $\chi_{2}$, the luminosity distance $D_{L}$, the merge time $t_{c}$, the right ascension $\alpha$ and declination $\delta$, the inclination angle $\iota$, the GW polarization angle $\psi$, and the phase of coalescence $\phi_{c}$.
When generating the BNS events, we adopt the population model proposed by Farrow et al. [74]. According to the spin magnitude, the model divides a BNS into a recycled NS and a nonrecycled (slow) one, for which the masses are labeled as $m_{\mathrm{r}}$ and $m_{\mathrm{s}}$, respectively. The distribution of $m_{\mathrm{r}}$ has two Gaussian components while $m_{\mathrm{s}}$ follows a uniform distribution,
$P(m_{\mathrm{r}})=\alpha\mathcal{N}(\mu_{1},\sigma_{1})+(1-\alpha)\mathcal{N}(\mu_{2},\sigma_{2})\,$, (3.1a)
$P(m_{\mathrm{s}})=\mathcal{U}(m_{\mathrm{s}}^{l},m_{\mathrm{s}}^{u})\,$, (3.1b)
[…]
We select a XG detector network consisting of two CE detectors and one ET detector, whose sensitivities are taken as CE-2 [54, 55] and ET-D [56, 57, 58], respectively. The two CE detectors are positioned at the current sites of the two LIGO detectors, while the ET detector is set at the current location of the Virgo detector with a triangular shape.
3.2 Implementation
According to the discussion in section 2.2, the inference of the Love-Q relation can be divided into two steps. In the auxiliary single-event inference, the variable parameters are
$\bm{\theta}=\{\mathcal{M},\eta,\Lambda_{1},\Lambda_{2},Q_{1},Q_{2},\chi_{1},\chi_{2},D_{L},t_{c},\alpha,\delta,\iota,\psi,\phi_{c}\}\,.$ (3.2)
In the parameter estimation, the priors of $\mathcal{M}$, $\eta$, $\chi_{1}$, $\chi_{2}$, $t_{c}$ and $\phi_{c}$ are uniform, and the prior of $D_{L}$ is such that the distribution is uniform in the co-moving volume. We set isotropic priors for the angle variables $\alpha,\delta,\iota,\psi$. For tidal and quadrupole moment parameters, we treat $\Lambda_{\mathrm{s}}$ and $Q_{\mathrm{s}}$ of the slow binary component as nuisance parameters, since the spin-induced quadrupole moment is poorly estimated for NSs when the spins are slow [39]. According to the arguments around Eq. (2.15), we choose uniform priors for $\Lambda_{1}$, $\Lambda_{2}$, $Q_{1}$ and $Q_{2}$ in the auxiliary inference. To calculate the posterior, we generate samples with the Bilby [81] implementation of the nessai sampler [82, 83, 84, 85, 86]. To calculate the integral in Eq. (2.9), we need the functional form of the quasi-likelihood, which is proportional to the distribution function of the posterior $p(\bm{\Lambda}_{i},\bm{Q}_{i}|d_{i},\varnothing)$. We follow Golomb and Talbot [70] and adopt the Gaussian mixture model to estimate the density of the posterior samples. In the second step, the priors of the hyperparameters, $\pi(\bm{H})$, are listed in table 1. For the conditional prior $\pi(\bm{\Lambda}_{i}|\bm{H})$, we choose the uniform distribution $\mathcal{U}(10,2000)$.
We simulate 1000 GW events as described in section 3.1. Though the hierarchical procedure avoids a direct high-dimensional inference for $\bm{H}$ and $\{\bm{\theta}_{1},\cdots,\bm{\theta}_{n}\}$ simultaneously, the computational cost still increases with the number of events $n$. Lackey and Wade [68] found that when constraining the EOS of NSs, several events with the highest signal-to-noise ratios (SNRs) contribute most of the information. Also, Yagi and Yunes [39] concluded that the spin-induced quadrupole moment cannot be well measured for NSs with low spins. Considering these findings, we first draw 100 sources with the highest SNRs, then further select 20 sources with the largest $|\chi_{\mathrm{r}}|$ among them. As we will show in section 4, the primary information for constraining the Love-Q relation comes from the 10 loudest GW events. We leave a more comprehensive study of the whole population for future work.
Table 1: The reference values and priors of the coefficients in different models of the Love-Q relation, with $j$ being the number of parameters in the polynomial model. Note that $j=5$ indicates the original Yagi-Yunes relation [43]. For $j=$ 2, 3 and 4, the values are the results of least squares regression performed on 1000 points uniformly picked in the logarithmic space from the Yagi-Yunes relation. Priors are chosen to be uniform distributions, and the ranges are the same for different $j$ to ensure a meaningful comparison.
Reference Values
$j$ $a_{j}$ $b_{j}$ $c_{j}$ $d_{j}$ $e_{j}$
5 $0.1940$ $0.0916$ $4.812\text{\times}{10}^{-2}$ $4.283\text{\times}{10}^{-3}$ $1.245\text{\times}{10}^{-4}$
4 $0.1290$ $0.1480$ $3.021\text{\times}{10}^{-2}$ $1.817\text{\times}{10}^{-3}$
3 0.0709-0.0709 $0.2775$ $3.220\text{\times}{10}^{-3}$
2 0.1457-0.1457 $0.3094$
Priors
$j$ $a_{j}$ $b_{j}$ $c_{j}$ $d_{j}$ $e_{j}$
5 $\mathcal{U}(-5.0,5.0)$ $\mathcal{U}(-1.0,1.0)$ $\mathcal{U}(-0.5,0.5)$ $\mathcal{U}(-0.1,0.1)$ $\mathcal{U}(-0.01,0.01)$
4 $\mathcal{U}(-5.0,5.0)$ $\mathcal{U}(-1.0,1.0)$ $\mathcal{U}(-0.5,0.5)$ $\mathcal{U}(-0.1,0.1)$
3 $\mathcal{U}(-5.0,5.0)$ $\mathcal{U}(-1.0,1.0)$ $\mathcal{U}(-0.5,0.5)$
2 $\mathcal{U}(-5.0,5.0)$ $\mathcal{U}(-1.0,1.0)$
4 Results and Discussions
In this section, we present the inference results for different parameterization models of the Love-Q relation. In section 4.1, we show the results of the linear model, Eq. (2.2), consisting of two parameters. In section 4.2, we present the results of the quartic polynomial model in Eq. (2.1), consisting of five parameters. The polynomial models in between, i.e., the quadratic and cubic polynomial models, are discussed in section 4.3.
4.1 The Linear Model
Figure 2: Posterior distributions of the hyperparameters ${\bm{H}}=\{a_{2},b_{2}\}$ in the linear fitting model. The contours refer to 50% and 90% credible regions, while the numbers above the histograms on the diagonal stand for the median and the central 90% credible interval of the marginalized distribution. We use blue and red colors to represent the results based on the loudest 20 and 10 events from the 1000 simulated events, respectively. The black lines represent the reference values in table 1. The grey lines on the diagonal represent the priors for comparison.
We first perform the inference for the linear model, which is also the model adopted in Ref. [64]. The posterior distribution of the hyperparameters $\{a_{2},b_{2}\}$ is shown in figure 2. Compared to the priors in table 1, the posteriors are significantly narrowed. As a comparison, we also mark the fitting values of the hyperparameters in a direct fit of the Yagi-Yunes Love-Q relation, and regard them as “true” values of the linear model for reference. These values fall within the 90% credible region of the posterior. The hierarchical Bayesian inference successfully recovers the Love-Q relation under the linear parameterization. We also investigate how the results depend on the number of events in the analysis. In the 20 events selected in section 3.2, we further select the loudest 10 events, perform the inference again, and show the results in figure 2. For both the joint and marginalized distributions, the widths of the credible regions in the 10-event inference are only slightly larger than those in the 20-event inference. This reveals that the loudest 10 events dominate the information in constraining the Love-Q relation, and including quieter events will not significantly change the results.
Figure 3: Recovered Love-Q relation from the posterior of the hyperparameters in the linear model. In the left panel, the Love-Q relation is inferred with the 20 loudest events from the 1000 simulated GW events. The gray points mark the median values of inferred $\Lambda$ and $Q$ for each event with $68\%$ errorbars. The blue solid line marks the median of the distribution of $Q$ as a function of $\Lambda$, accompanied by the $50\%$ and $90\%$ credible intervals in shaded regions. The red dashed line represents the maximum-posterior Love-Q relation. For comparison, we plot the original Yagi-Yunes Love-Q relation [43] in orange. The right panel shows how the $90\%$ credible region of the recovered Love-Q relation depends on the number of events, marked with different colors.
In figure 3, we show the recovered Love-Q relation according to the posterior samples. For each $\Lambda$, every sample in the posterior of the hyperparameters corresponds to a $Q$ value. For a fixed $\Lambda$ value, we find the credible intervals of $Q$, and then vary $\Lambda$ continuously to form credible regions. In the left panel, we show the results of the 20-event inference. Similar to figure 2, the Yagi-Yunes Love-Q relation is covered by the 90% credible region. In the right panel, we select the loudest 5, 10, 15 and 20 events from the 20 events to test how the recovered Love-Q relation depends on the number of events. We find that the widths of the 90% credible regions are almost the same for the inferences from 10, 15 and 20 events. This is consistent with the posteriors in figure 2, and again indicates that the loudest 10 events are almost sufficient to constrain the Love-Q relation. Similar phenomena were also found in previous studies [68, 87, 88, 89, 90, 72], where the recovered $\Lambda$-$m$ relation is dominated by the several loudest events.
4.2 The Quartic Polynomial Model
Figure 4: The posterior of the hyperparameters and the recovered Love-Q relation in the quartic polynomial model, where all 20 loudest events are included in the inference. Figure settings of two panels are respectively similar to those in figure 2 and the left panel of figure 3.
For the quartic polynomial model, the Love-Q relation is fitted with five parameters shown in Eq. (2.1), i.e., $\{a_{5},b_{5},c_{5},d_{5},e_{5}\}$. This is the original model proposed by Yagi and Yunes [39]. We summarize the posterior and the recovered Love-Q relation in figure 4, where all 20 events are included in the inference. In the left panel, though the true values (the values in the Yagi-Yunes Love-Q relation [39]) are almost centered in the distribution, the posteriors are much wider than those in the linear case, reaching the prior boundaries. This indicates that the $\Lambda$ and $Q$ measurements from these events are not informative enough to well constrain all the five parameters. In other words, the observation precision is not high enough to capture the higher-order terms introduced by the additional three parameters, $c_{5},d_{5}$, and $e_{5}$. This is also reflected in the strong correlations between the three parameters shown in the left panel of figure 4. In the right panel of figure 4, the 90% credible regions of the recovered Love-Q relation are also wider than that in the linear case, especially for large $\Lambda$ where the higher-order terms become more important. For $\Lambda\sim 400$, the widths of the 90% credible region between the two models are similar, which is consistent with the fact that most of the simulated events gather around this region.
4.3 The Quadratic and Cubic Polynomial Models
Figure 5: Similar to figure 4, while for the quadratic and cubic polynomial models.
As shown in the previous two subsections, the observations from 20 simulated GW events can well constrain the two parameters in the linear model, while cannot effectively constrain the five parameters in the quartic polynomial model. The above two models were adopted in previous studies [39, 64], and here we further test the models in between them and see how the constraints change with the number of hyperparameters. We perform the inference for the quadratic and cubic polynomial models, which contain three parameters, $\{a_{3},b_{3},c_{3}\}$, and four parameters, $\{a_{4},b_{4},c_{4},d_{4}\}$, respectively. Similar to the linear model, the reference values of these parameters are obtained by directly fitting the Yagi-Yunes Love-Q relation and are listed in table 1. The priors of the hyperparameters are kept the same as their counterparts in the quartic polynomial model, which are also listed in table 1.
We summarize the posterior distributions of the hyperparameters in the top row of figure 5. In the quadratic model, strong degeneracy arises among the three parameters. The posterior of $b_{3}$ is almost the same as its prior. For $a_{3}$ and $c_{3}$, though the posteriors are narrower than the priors, their marginal distributions have wide and flat plateau. In the cubic model, strong degeneracy exists between $c_{4}$ and $d_{4}$, and the posterior of $b_{4}$ is similar to its prior. In general, posterior with large correlations and prior-like marginal distributions indicates redundant parameters in the model. Therefore, we conclude that the linear model is accurate enough in constraining the Love-Q relation with XG GW observations. This is consistent with the argument in Ref. [64] based on qualitative analysis, whereas in this work we present a quantitative demonstration.
Same as before, we plot recovered Love-Q relations in the lower panels of figure 5. We find that the widths of the 90% credible regions around $\Lambda\sim 400$ are similar for all of four models. While for $\Lambda$ away from this region ($\Lambda\lesssim 200$ or $\Lambda\gtrsim 500$), the widths increase with the number of parameters. This may have resulted from the increased model complexity and the lack of data points in these regions.
5 Testing Modified Gravity: Dynamical Chern-Simons Gravity
Figure 6: Comparison of the recovered Love-Q relation in GR and that are predicted by the dCS gravity with coupling constant $\bar{\xi}$ fixed. Similar to figure 3, the blue lines indicate the median of the recovered Love-Q relation, while the shaded regions represent the $50\%$ and $90\%$ credible intervals. For the coupling constant $\bar{\xi}$, we take three values and plot the corresponding Love-Q relations.
In this section, we investigate the potential of testing gravity theories with the inferred Love-Q relation from the GW observations. The I-Love-Q test can be powerful when significant difference in I-Love-Q relation between GR and modified gravity theories exists [1], for example, in some parity-violating theories [43, 45]. Here we take the dynamical Chern-Simons (dCS) gravity [91, 92, 93] for illustration.
The dCS gravity is well-motivated from heterotic superstring theory, loop quantum gravity, and effective field theories of inflation [91, 92, 93]. The dCS gravity introduces parity violation and quadratic curvature terms in the action [93, 44],
$S=\int\mathrm{d}^{4}x\sqrt{-g}\left[\kappa_{g}\mathcal{R}+\frac{\alpha}{4}\mathcal{\vartheta}\mathcal{R}_{\nu\mu\rho\sigma}{}^{*}\mathcal{R}^{\mu\nu\rho\sigma}-\frac{\beta}{2}\nabla_{\mu}\mathcal{\vartheta}\nabla^{\mu}\mathcal{\vartheta}+\mathcal{L}_{\mathrm{mat}}\right]\,$, (5.1)
where $g$ is the determinant of the metric, $\kappa_{g}=1/16\pi$, $\mathcal{R}$ is the Ricci scalar, $\mathcal{R}_{\nu\mu\rho\sigma}$ and $^{*}\mathcal{R}^{\mu\nu\rho\sigma}$ denote the Riemann tensor and its dual, $\mathcal{L}_{\mathrm{mat}}$ is the matter Lagrangian density, $\alpha$ and $\beta$ are the coupling constants in the dCS gravity and the potential for the pseudo-scalar field $\mathcal{\vartheta}$ is omitted. $\mathcal{\vartheta}$ and $\beta$ is taken to be dimensionless, so $\alpha$ has the dimension of length square, and the quantity $\xi_{\mathrm{CS}}^{1/4}\equiv[\alpha^{2}/(\kappa\beta)]^{1/4}$ can be explained as the characteristic lengthscale of the dCS gravity [94, 95]. Current Solar System observations have constrained it to $\xi_{\mathrm{CS}}^{1/4}<\mathcal{O}(10^{8})$ km [96, 95].
The dCS gravity predicts Love-Q relations that deviate from the one in GR [38, 39, 44], allowing us to test the dCS gravity with the inferred Love-Q relation. For a Love-Q test, Yagi et al. [97] have obtained the dCS correction to the NS quadrupole moment, and Yagi et al. [98] indicates that, regarding the dCS gravity as an effective theory, the tidal deformability is the same as in GR at leading order in the small coupling approximation, $\zeta\equiv\xi_{\mathrm{CS}}m^{2}/R^{6}\ll 1$. Refs. [43, 97, 44] have discussed the Love-Q relation in the dCS gravity and found that the relation becomes EOS-sensitive with $\xi_{\mathrm{CS}}$ or $\zeta$ fixed. However, with $\bar{\xi}\equiv\xi_{\mathrm{CS}}/m^{4}$ fixed, the Love-Q relation remains universal and the variation with respect to EOS is of $\mathcal{O}(1\%)$. Gupta et al. [44] parameterized the relation between the dCS correction to $Q$ (denoted as $Q_{\mathrm{CS}}$) and the tidal deformability $\Lambda$ as
$\ln(Q_{\mathrm{CS}}/\bar{\xi})=a+b\ln\Lambda+c\ln^{2}\Lambda\,$, (5.2)
where the fitting coefficients are a=3.443,b=0.550a=-3.443,b=-0.550, and c=0.023c=-0.023.
We compare the inferred constraints of the Love-Q relation using the linear model from figure 3 and Love-Q relations (5.2) with $\bar{\xi}$ fixed in the dCS gravity. From figure 6 we conclude that our Love-Q test of the dCS gravity can constrain the coupling constant to $\bar{\xi}\lesssim 10^{3}$. Substituting a typical mass $m=1.4\,\mathrm{M_{\odot}}$ and a typical radius $R=10$ km for NSs, we have constraints for the other two coupling constants, $\xi_{\mathrm{CS}}^{1/4}\lesssim 10$ km and $\zeta\lesssim 0.1$, which are in agreement with the results given by Refs. [38, 39].
6 Conclusion
In this work, we investigate the prospects of inferring the Love-Q relation of NSs with future ground-based GW observations. Extending the inference procedure by Samajdar and Dietrich [64], we adopt the hierarchical Bayesian framework to effectively combine the information from multiple GW events, extending the linear fitting model in Ref. [64]. The hierarchical Bayesian framework separates the inferences into two steps, the auxiliary single-event inference and the hyperparameter inference, which avoids a direct high-dimensional inference for all the parameters simultaneously while still takes into account degeneracy and non-Gaussianity in the single-event parameters. We also discuss the impact of the event number on the constraints, and find that the loudest 10 events dominate the information in constraining the Love-Q relation. Similar phenomena were also found in studies of constraining the EOSs with GW observations [68, 87, 88, 89, 90, 72].
To our knowledge, we are the first to conduct a systematic study on different parameterization models when inferring the Love-Q relation with GWs. Considering the quartic polynomial model proposed by Yagi and Yunes [39] and the linear model adopted by Samajdar and Dietrich [64] as two ends, we conduct the inference with four polynomial models from linear to quartic terms. As shown in section 4, we quantitatively demonstrate that the linear model is accurate enough to describe the Love-Q relation in the inference in the CE and ET era. As the number of hyperparameters that parametrize the Love-Q relation increases, more significant degeneracy or poorly constrained posteriors appear, while the recovered Love-Q relations only keep similar in the region where most data points gather.
We also test the potential of using the inferred Love-Q relation to constrain modified gravity theories in section 5. Taking the dCS gravity as an example, we find that the inferred Love-Q relation can place a constraint on the dCS characteristic lengthscale to about $\xi_{\mathrm{CS}}^{1/4}\lesssim 10$ km, which is seven orders of magnitude tighter than that from the current Solar System observations [96, 95], and is consistent with previous predictions with GWs [38, 39]. This highlights the power of inferring the Love-Q relation from GWs in testing gravity theories in the strong-field regime.
There can be several extensions to this study. Firstly, we assume BNS systems with aligned spins, while realistic BNS systems may have tilted spins and spin precession, which leads to additional contribution in the GW waveform [99, 100, 101]. Additionally, in the XG era, the GW signals are possibly overlapping with each other, complicating the data analysis [102, 63, 103, 104, 105]. Moreover, the Yagi-Yunes universal relations were originally found and discussed for a single NS. For BNS systems, these universal relations may hold when the two NSs have similar masses, but still require a careful inspection [1, 106]. The octupole or higher-order multipole moments of NSs also contribute to the GW waveform and can be universally related to the quadrupole moment [43, 18]. We leave the exploration of simultaneous or independent inference of more universal relations to future studies.
Acknowledgments
This work was supported by the Beijing Natural Science Foundation (QY25102, 1242018), the National Natural Science Foundation of China (123B2043, 12573042), the National SKA Program of China (2020SKA0120300), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-Performance Computing Platform of Peking University.

### v2

@@
 The existence of such universal relations, which now include many others [59, 111, 64, 70, 109, 39, 48], may also imply a “no-hair-theorem-like” behavior for NSs, bringing us a new insight into fundamental physics.
-Future next-generation (XG) ground-based GW detectors, including the Cosmic Explorer (CE) [54, 55] and the Einstein Telescope (ET) [56, 57, 58, 59], are expected to detect many more GW signals, up to about $10^{5}$$10^{6}$ events per year [60, 61, 62, 63], thanks to their increased sensitivity and lower cutoff frequencies.
+Future next-generation (XG) ground-based GW detectors, including the Cosmic Explorer (CE) [76, 77] and the Einstein Telescope (ET) [74, 46, 82, 3], are expected to detect many more GW signals, up to about $10^{5}$$10^{6}$ events per year for BNS coalescences [8, 83, 52, 81], thanks to their increased sensitivity and lower cutoff frequencies.
 These high-precision GW observations allow us to treat $\Lambda$ and $Q$ as independent parameters in the waveform model, to be measured directly from GWs.
@@
 2.1 Polynomial Models of Love-Q Relation
-Yagi and Yunes [38, 39, 43] fit the Love-Q relation with a quadratic polynomial model as,
+Yagi and Yunes [108, 107, 110] fit the Love-Q relation with a quartic polynomial model as,
 $\ln Q_{5}=a_{5}+b_{5}\ln\Lambda+c_{5}\ln^{2}\Lambda+d_{5}\ln^{3}\Lambda+e_{5}\ln^{4}\Lambda\,$, (2.1)
@@
 These hyperparameters do not enter the waveform directly, while can be inferred from a collection of measurements of single-event parameters.
-Similarly, the EOS parameters determine the relation between $\Lambda$ and $m$, as well as between $Q$ and $m$, can be regarded as hyperparameters and analyzed in the hierarchical Bayesian framework [65, 66, 67, 68, 69, 70, 71, 72].
+The EOS parameters, which determine the $\Lambda$-$m$ and $Q$-$m$ relations, can also be regarded as hyperparameters and analyzed in the hierarchical Bayesian framework [62, 63, 10, 56, 61, 40, 9, 95].
 In this work, we adopt this framework to the inference of the Love-Q relation.
@@
 $\pi(\bm{H},\bm{\theta}_{1},\cdots,\bm{\theta}_{n})=\pi(\bm{H})\prod_{i=1}^{n}\pi(\bm{\theta}_{i}|\bm{H})\,.$ (2.6)
-Since we are inferring the Love-Q relation, the tidal deformabilities, $\bm{\Lambda}_{i}=\{\Lambda_{1i},\Lambda_{2i}\}$, and quadrupole moments, $\bm{Q}_{i}=\{Q_{1i},Q_{2i}\}$, of the two NSs of the $i\text{-th}$ event are treated as independent parameters in $\bm{\theta}_{i}$.
+We treat the tidal deformabilities, $\bm{\Lambda}_{i}=\{\Lambda_{1i},\Lambda_{2i}\}$, and quadrupole moments, $\bm{Q}_{i}=\{Q_{1i},Q_{2i}\}$, of the two NSs of the $i\text{-th}$ event as independent parameters in $\bm{\theta}_{i}$ and select flat priors for them.
+The impacts of priors are relatively small, since only the loudest events with high SNRs are analyzed.
 Note that the Love-Q relation is insensitive to EOS, the priors of mass parameters, $m_{1}$ and $m_{2}$, are chosen to be independent of $\bm{\Lambda}_{i},\bm{Q}_{i}$ and $\bm{H}$.
@@
 In our simulation, we adopt the IMRPhenomXAS_NRTidalv3 waveform model [2], which includes tidal amplitude corrections as well as spin-induced quadrupole moment terms up to 3.5 PN with aligned spins.
-The single-event parameters $\bm{\theta}$ include the binary masses $m_{1}$ and $m_{2}$, the dimensionless tidal deformabilities $\Lambda_{1}$ and $\Lambda_{2}$, spin-induced quadrupole moments $Q_{1}$ and $Q_{2}$, the dimensionless spins $\chi_{1}$ and $\chi_{2}$, the luminosity distance $D_{L}$, the merge time $t_{c}$, the right ascension $\alpha$ and declination $\delta$, the inclination angle $\iota$, the GW polarization angle $\psi$, and the phase of coalescence $\phi_{c}$.
+The single-event parameters $\bm{\theta}$ include the binary masses $m_{1}$ and $m_{2}$, the dimensionless tidal deformabilities $\Lambda_{1}$ and $\Lambda_{2}$, spin-induced quadrupole moments $Q_{1}$ and $Q_{2}$, the dimensionless spins $\chi_{1}$ and $\chi_{2}$, the luminosity distance $D_{L}$, the coalescence time $t_{c}$, the right ascension $\alpha$ and declination $\delta$, the inclination angle $\iota$, the GW polarization angle $\psi$, and the phase of coalescence $\phi_{c}$.
 When generating the BNS events, we adopt the population model proposed by Farrow et al.
@@
 $P(m_{\mathrm{s}})=\mathcal{U}(m_{\mathrm{s}}^{l},m_{\mathrm{s}}^{u})\,$, (3.1b)
+where $\alpha=0.68$, $\mu_{1}=1.34\,\mathrm{M}_{\odot}$, $\sigma_{1}=0.02\,\mathrm{M}_{\odot}$, $\mu_{2}=1.47\,\mathrm{M}_{\odot}$, $\sigma_{2}=0.15\,\mathrm{M}_{\odot}$, and $m_{\mathrm{s}}^{l}=1.14\,\mathrm{M}_{\odot}$, $m_{\mathrm{s}}^{u}=1.46\,\mathrm{M}_{\odot}$.
+The NS with a larger mass is labeled as the primary star with mass $m_{1}$ and the other as the secondary star with mass $m_{2}$.
+The tidal deformability and quadrupole moments of the binary are calculated from the stellar mass assuming the APR4 EOS with methods described in Refs.
+[107, 19] and the slow rotation approximation is used.
+For arbitary spin, Refs.
+[29, 71, 103, 28] have developed new relations insensitive to certain set of EOSs with fitting coefficients depending on the spin parameter.
+Ref.
+[22] further suggested introducing the fitting coefficients as a function of both spin and radius, extending the universality to various EOSs.
+In these cases, our framework will still be useful with appropriate model parameterizations and modifications.
 […]
@@
 The two CE detectors are positioned at the current sites of the two LIGO detectors, while the ET detector is set at the current location of the Virgo detector with a triangular shape.
+The current locations of LIGO and Virgo are chosen as an illustrated configuration for XG detectors.
 3.2 Implementation
@@
 Considering these findings, we first draw 100 sources with the highest SNRs, then further select 20 sources with the largest $|\chi_{\mathrm{r}}|$ among them.
-As we will show in section 4, the primary information for constraining the Love-Q relation comes from the 10 loudest GW events.
-We leave a more comprehensive study of the whole population for future work.
+The SNR values of these selected events range from about 1500 to 2500.
+Here we omit the selection effects since different $\Lambda$ and $Q$ values do not lead to a significant change of SNR and thus the detection probability.
+As we will show in section 4, the primary information for constraining the Love-Q relation comes from the loudest GW events.
+In practice, selecting the top 20 events is enough to capture the main information for constraining the Love-Q relation.
+We present the results of changing the number of selected events in appendix A.
+We leave a more comprehensive study of the whole population considering selection effects for future work.
 Table 1: The reference values and priors of the coefficients in different models of the Love-Q relation, with $j$ being the number of parameters in the polynomial model.
@@
 The polynomial models in between, i.e., the quadratic and cubic polynomial models, are discussed in section 4.3.
+For other parameterizations beyond polynomial models, our methodology is in principle applicable to them as well.
 4.1 The Linear Model
@@
 We use blue and red colors to represent the results based on the loudest 20 and 10 events from the 1000 simulated events, respectively.
-The black lines represent the reference values in table 1.
+The orange bands represent the reference values with fitting uncertainties accounted for.
 The grey lines on the diagonal represent the priors for comparison.
@@
 These values fall within the 90% credible region of the posterior.
+Given that the Love-Q relation is quasi-universal with relative differences of $\sim 1\%$ for various EOSs, the fitting uncertainties of the hyperparameters are also considered in the reference values.
+To check the robustness of our conclusions, we also repeat the inference assuming the SLy EOS; which are summarized in appendix B.
 The hierarchical Bayesian inference successfully recovers the Love-Q relation under the linear parameterization.
@@
 where $g$ is the determinant of the metric, $\kappa_{g}=1/16\pi$, $\mathcal{R}$ is the Ricci scalar, $\mathcal{R}_{\nu\mu\rho\sigma}$ and $^{*}\mathcal{R}^{\mu\nu\rho\sigma}$ denote the Riemann tensor and its dual, $\mathcal{L}_{\mathrm{mat}}$ is the matter Lagrangian density, $\alpha$ and $\beta$ are the coupling constants in the dCS gravity and the potential for the pseudo-scalar field $\mathcal{\vartheta}$ is omitted.
-$\mathcal{\vartheta}$ and $\beta$ is taken to be dimensionless, so $\alpha$ has the dimension of length square, and the quantity $\xi_{\mathrm{CS}}^{1/4}\equiv[\alpha^{2}/(\kappa\beta)]^{1/4}$ can be explained as the characteristic lengthscale of the dCS gravity [94, 95].
+$\mathcal{\vartheta}$ and $\beta$ are taken to be dimensionless, so $\alpha$ has the dimension of length squared, and the quantity $\xi_{\mathrm{CS}}^{1/4}\equiv[\alpha^{2}/(\kappa\beta)]^{1/4}$ can be explained as the characteristic lengthscale of the dCS gravity [112, 106].
 Current Solar System observations have constrained it to $\xi_{\mathrm{CS}}^{1/4}<\mathcal{O}(10^{8})$ km [16, 106].
@@
 Firstly, we assume BNS systems with aligned spins, while realistic BNS systems may have tilted spins and spin precession, which leads to additional contribution in the GW waveform [102, 75, 37].
+The inaccuracy of waveform itself can also introduce a theoretical error to the posteriors of single event parameters, especially for events with high SNRs [25].
 Additionally, in the XG era, the GW signals are possibly overlapping with each other, complicating the data analysis [72, 81, 97, 51, 96].
-Moreover, the Yagi-Yunes universal relations were originally found and discussed for a single NS.
-For BNS systems, these universal relations may hold when the two NSs have similar masses, but still require a careful inspection [1, 106].
+Moreover, the Yagi-Yunes universal relations were originally found and discussed for a single NS under the slow-rotation assumption.
+For NSs in BNS systems or with high spins, these universal relations may break down and new relations could emerge.
+Our methodology in principle could still work when new universal relations do exist, but a careful inspection is required [85, 78, 29, 71, 22, 103, 28, 101].
 The octupole or higher-order multipole moments of NSs also contribute to the GW waveform and can be universally related to the quadrupole moment [110, 2].
@@
 Acknowledgments
+We thank the anonymous referee for helpful comments.
 This work was supported by the Beijing Natural Science Foundation (QY25102, 1242018), the National Natural Science Foundation of China (123B2043, 12573042), the National SKA Program of China (2020SKA0120300), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-Performance Computing Platform of Peking University.
+Appendix A Dependence on the Number of Events
+In the main text, we have selected the loudest 20 events from the 1000 simulated events in the inference.
+Here we vary the number of events $N_{\text{events}}$ to see how the results change.
+As shown in figure 7, for a small $N_{\text{events}}$ (like $N=5$), there is a noticeable offset between the median estimate and the reference value of the hyperparameters.
+However, as $N_{\text{events}}$ increases, both the median and the credible intervals shrink towards the reference values.
+to the reference values.
+Due to the decreasing SNRs of the latter added samples, their contribution to the inference becomes less and less significant, and the results keep almost unchanged.
+The choice of 20 loudest events can be justified with figure 8, where we observe that the constraint to Love-Q relation at $\Lambda\sim 350$ (the “waist” part in figure 3 of the manuscript) becomes stable when $N_{\text{events}}$ is larger than 20.
+Similar results are also found in previous studies of constraining the EOSs with GWs, where the several loudest events dominate the information [56].
+Figure 7: The dependency of the 5th, 50th and 95th percentiles of the hyperparameter posterior samples on the number of events.
+The orange dashed line represents the reference values, i.e.
+the linear fitting values obtained in section 2.1.
+Figure 8: The dependency of 5th and 95th percentiles of $Q$ samples computed at $\Lambda=350$ (where most of the data points gather) on the number of events.
+The orange dashed line represents $Q$ calculated from the Yagi-Yunes relation.
+Appendix B Results for SLy
+Different assumptions of EOS provide different combinations of $\Lambda$ and $Q$, which still follow the universal relation.
+As a supplement, we have repeated the inference of linear model for SLy, another EOS consistent with GW170817 [4, 5, 6].
+The results are presented and compared with those of APR4 in figure 9 and figure 10.
+As shown in figure 9, the uncertainties of fitting coefficients brought by multiple EOSs are much smaller compared to those from the hierarchical inference.
+Thus the change of EOS would not introduce a significant change to the width of constraints, while the median may probably change with EOSs.
+And in figure 10.
+The median and maximum posterior estimations change, but the width of 90% credible region remains insensitive to the choice of EOS.
+Figure 9: Corner plots for APR4 and SLy EOSs.
+Figure 10: Constraints to Love-Q relation in a linear model for APR4 and SLy.